# On Inverse Problems, Parameter Estimation, and Domain Generalization

### v1

@@ 6 Examples
Definition 13 (Surjective mapping)
@@ 6.1 Deterministic mappings vs stochatic mappings
Traditionally, most signal restoration algorithms were deterministic, i.e., the algorithm provides one reconstruction solution for a given degraded input. Mathematically speaking, assume $y=g(x)$ such that $g(\cdot)$ is a deterministic function, and $f(\cdot)$ is a surjective function. Therefore, $p(X=f(y)|Y=y)=1$, or equivalently, $p_{X|Y}(x|y)=\delta(x-f(y))$. Given a particular input $Y=y$, a restoration algorithm produces an estimate $\hat{X}$ according to some distribution $p_{\hat{X}|Y}(\cdot|y)$ such that the estimate $\hat{X}$ is statistically independent of $X$ given $Y$ For deterministic algorithms, $p_{\hat{X}|Y}(\cdot|y)$ is a delta function for every $y$, while for stochastic algorithms it is a non-degenerate distribution.
Recall,
$H(\theta|X)=\sum_{x}p(x)H(\theta|X=x)=-\sum_{x}p(x)\sum_{\theta}p(\theta|x)\log p(\theta|x)=-\sum_{x}\sum_{\theta}p(x,\theta)\log p(\theta|x).$ (49)
Assuming $\theta-Y-X$ forms a Markov chain, $P(\theta,X)=P(\theta,X|Y=T(X))$, and $H(\theta|X)\leq H(\theta|Y)$, with equality if $g(\cdot)$ is a bijective function. Thus, in case that $\mathcal{F}:\mathcal{Y}\rightarrow\mathcal{X}$ is a surjective mapping fully captured by the estimator of $\hat{X}$, i.e., $p_{X|Y}(\cdot|y)=p_{\hat{X}|Y}(\cdot|y)$, as stated in theorems 10-11 an algorithm that is able to extract the source signal conditional distribution would yield a parameter estimation result that is as accurate as the actual source and as the measurement. The simplest example would be $Y=aX$, where $a>0$ is a constant.
In the case that $y=g(x)$ where $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ is a non-invertible deterministic degradation such that $p_{Y|X}(y|x)=\delta(y-g(x))$, Ohayon et al. (2023a) proved that when a restoration algorithm enforces both perfect perceptual quality , i.e., $p_{\hat{X}}=p_{X}$ and perfect data fidelity, that is $y=g(\hat{x})$ (equivalent to $p_{Y|\hat{X}}=p_{Y|X}$), it must be a sampler from the posterior distribution $p_{X|Y}$, and is thus required to be stochastic. In other words, these are ill-posed inverse problems in which $X$ cannot be retrieved from $Y$ with zero error ($p_{X|Y}(\cdot|y)$ is not a delta function for almost every $y$). Problems such as image colorization, inpainting, single-image super-resolution, JPEG-deblocking, and more all follow this assumed structure. Thus, in these cases, as indicated in Theorems 10-12, in order to maintain the same level of accuracy for parameter estimation after the reconstruction, there is need for prior knowledge in the restoration process, such as an oracle of $\theta$, and/or a restoration process that does not depend on $\theta$.
@@ 6.2 When $\hat{\theta}_{Y}$ and $\hat{\theta}_{\hat{X}}$ coincide
Let us assume, $Y=X+Z$, where $Z\sim\mathcal{N}(0,\sigma_{\mathrm{n}}^{2})$ and $X$ are independent. Hence $Z$ and $\theta$ are independent. That is, $\theta-X-Y$ forms a Markov chain. As known, generally speaking, in this case, averaging examples of $Y$ yields $X$, that is, $EY=X$.
@@
Now, assume $Y=X+Z$, then $Y\sim\mathcal{N}(\mu,\sigma^{2}_{\mathrm{y}})$ such that $\sigma^{2}_{\mathrm{y}}=\sigma^{2}_{\mathrm{x}}+\sigma^{2}_{\mathrm{z}}$. Assuming that $\theta=\mu\sim\mathcal{N}(\mu_{\mathrm{m}},\sigma^{2}_{\mathrm{m}})$, we have $I(\theta,X)=\frac{1}{2}\mathrm{log}(1+\frac{\sigma^{2}_{\mathrm{m}}}{\sigma^{2}_{\mathrm{x}}-\sigma^{2}_{\mathrm{m}}})$, and $I(\theta,Y)$=12log($\frac{1}{2}\mathrm{log}(1+\frac{\sigma^{2}_{\mathrm{m}}}{\sigma^{2}_{\mathrm{y}}-\sigma^{2}_{\mathrm{m}}})$. As well as $J^{-1}_{m,X}(\theta)=\frac{\sigma_{\mathrm{x}}^{2}}{m}$ and $J^{-1}_{m,Y}(\theta)=\frac{\sigma_{\mathrm{y}}^{2}}{m}$. At first glance, this may seem as a contradiction, since, obviously $I(\theta,X)>I(\theta,Y)$. Yet, provided that we would be able to recover $\hat{x}$, such that $p_{\hat{X}|\theta}(x|\theta)=p_{X|\theta}(x|\theta)$, which may be approximately possible in practice for some cases, under sufficiently low noise levels, we should be able to achieve the minimal CRB. Recall, that for a memory-less source we have
$I(\mathbf{x},\mathbf{y})\geq\sum^{m}_{i=1}I(\mathbf{x}_{i},\mathbf{y}_{i}),$ (51)
@@
An optimal linear estimator for $X$ in the above example would be $\hat{x}=ay+b$ such that $a=\frac{\sigma_{x}}{\sigma_{y}}$, $b=E[Y](1-\frac{\sigma_{\mathrm{x}}}{\sigma_{\mathrm{y}}})$. In that case, in accordance with Theorem 11, clearly $P_{\hat{X}}(x)=P_{X}(x)$, yet obviously $\forall\theta\ :\ \hat{x}\neq x$.
@@ A Gaussian encoder (variance preserving) at time step $t$,
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\sqrt{\alpha_{t}}\mathbf{x}_{t-1},(1-\sqrt{\alpha_{t}})\mathbf{I}).$ (52)
We are interested in learning conditionals $p_{\phi}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, so that we can simulate new data. Intuitively, the forward process gradually adds noise to the observation $\mathbf{x}_{0}$, whereas the generative process gradually denoises a noisy observation. If we assume $\theta=\mathbf{x}_{0}$ is the clean image, and $\mathbf{x}=\mathbf{x}_{t}$ such that $0<t<T$ and $\mathbf{y}=\mathbf{x}_{T}$ is the noise image. The generative process approximates an intractable reverse process $p_{\phi}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$. The training loss (Ho et al., 2020) uses KL divergence to directly compare $p_{\phi}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ against forward process posteriors, which are tractable when conditioned on $\mathbf{x}_{0}$.
@@ 6.3 Sparse Inversion
$=\arg\min_{\mathbf{x}\in\mathbb{R}^{N}}\quad\log{p_{Y|X}(\mathbf{y}|\mathbf{x})}-\log{p_{X}(\mathbf{x})}.$ (53)
@@
$\hat{\mathbf{x}}=\arg\min_{\mathbf{x}\in\mathbb{R}^{N}}\quad\frac{1}{2\sigma_{\mathrm{z}}^{2}}\|\mathbf{y-Hx}\|_{2}^{2}-\lambda\|\mathbf{x}\|_{1},$ (54)
which is considered in the literature as the well known sparse coding $\ell_{1}$ optimization problem (see (Elad, 2010), Chapter 6). The first term is (54) is the data fidelity term, and the second term is the regularization term. The sparse coding recovery estimation is an extensively studied problem (Fernandez-Granda, 2013; Candès and Fernandez-Granda, 2013; Pereg and Cohen, 2017; Pereg et al., 2021). Pereg et al. (2019) proved that the bound on the estimation of $\mathbf{x}$ under sufficient separation between adjacent non-zero values (spikes) depends on the noise level. In the absence of noise, the MAP estimation should be perfect. However, note that solving (54) accurately requires the prior knowledge of $\lambda$.
@@
$s^{m}(\lambda)=\frac{\partial}{\partial\lambda}\log p^{m}_{\lambda}(\mathbf{x})=\frac{\partial}{\partial\lambda}\big{(}m\log(\lambda)-\lambda\sum_{i=1}^{m}\|\mathbf{x}_{i}\|_{1}\big{)}.$ (55)
@@
$J^{m}_{X}(\lambda)=mJ_{X}(\lambda)=-mE\Bigg{\{}\Bigg{(}\frac{\partial^{2}}{\partial\lambda^{2}}\log p_{\lambda}(\mathbf{x})\Bigg{)}^{2}\Bigg{|}\lambda\Bigg{\}}=\frac{m}{\lambda^{2}},$ (56)
In other words, hypothetically, if we were to first perfectly denoise the signal such that $\|\hat{\mathbf{x}}\|_{1}=\|\mathbf{x}\|_{1}$, assuming that our denoising mechanism is errorless, we have $\lambda^{-1}_{\hat{X}}=\|\hat{\mathbf{x}}\|_{1}$ ($m=1$). That said, by Theorem 10 we know that even in that case $\Delta_{Y}=\Delta_{\hat{X}}$. In other words, if we were to estimate $\lambda$ directly from the measurement, then our expected bound on the parameter estimation error is the same. If we solve the MAP for noisy data the upper-bound for $\hat{\mathbf{x}}$ is given by eq. (68),(72) in Appendix B, and it cannot be guaranteed that $\lambda_{\hat{X}}=\lambda$, namely, $\Delta_{Y}\leq\Delta_{\hat{X}}$. Interestingly, in the noisy case, denoising the measure will not improve the following expected estimation error.
@@
Let us assume $x(t)=\delta(t-\theta)$, in other words $p_{X|\theta}$ is a delta function. Assuming $y=h*x+w$, if we were able to recover $p_{\hat{X}|\theta}(x|\theta)=p_{X|\theta}(x|\theta)$, then we know $\theta_{\hat{X}}=\theta$. According to Theorem 10, in this case we have $I(\theta;X)=I(\theta;Y)=I(\theta;\hat{X})$ and $\Delta_{Y}=\Delta_{\hat{X}}$. In other words, for the sake of recovering $\theta$ we gain nothing from denoising $y$. It could of course be valuable to us to denoise $y$ for perceptual visual evaluation, and in some cases practically more convenient or possible. Yet, in terms of the accuracy of the parameter recovery we do not gain from first denoising $y$.
Note that a similar formulation can be applied to 2D cases where $\theta$ is the support of any object we wish to detect in an image. The question of bounding the error in this case-study remains an open question that we leave for future work.
7 Discussion &\&& Conclusions
When working on inverse problems in different domains, we tend to assume that the inversion process cleans the data and extracts the relevant information. We often forget that the process could also be inherently biased and inherently inclined to lose information. This study broadly analyzes parameter estimation and domain shift in inverse problems. We introduced general definitions for perfect perceptual estimators in the context of parameter estimation in continuous and discrete settings, and presented theoretical bounds for reliable prediction. Observing that domain randomization and data augmentation methods attempt the build models that generalize over several distributions, we highlight a profound vulnerability of this approach in the Double Meaning Theorem (Theorem 3), which often leads to mode collapse in signal restoration tasks. Our theoretical findings are confirmed in experimental examples. Future work can analyze other specific distributions depending on parameter estimation, such as: Bernoulli distribution, and mixture models e.g., product of experts (Hinton, 1999), experimentally investigate other types of domain changes and adversarial attacks in different applications, and develop coping strategies. Our future work will investigate how bias during training affects the classification decision, and the effect of number of measurements, and sample size.
Acknowledgments and Disclosure of Funding
The author wishes to thank George Barbastathis for helpful discussions.
@@ Appendix A Information-Theory Preliminaries
A.1 Efficiency
@@
$V(\hat{\theta})=\frac{J^{-1}(\theta)}{\mathrm{Var}(\hat{\theta})},$ (57)
@@
$\mathrm{Var}(\hat{\theta})=E(\theta-\hat{\theta})^{2}.$ (58)
@@ A.2 Consistency
A.3 Minimal Sufficient Statistic
Definition 14 (sufficient statistic (Cover and Thomas, 2006))

### v2

@@ 6 Examples
Definition 6.1 (Surjective mapping).
@@ 6.1 Deterministic mappings vs stochatic mappings
Traditionally, most signal restoration algorithms were deterministic, i.e., the algorithm provides one reconstruction solution for a given degraded input. Mathematically speaking, assume $y=g(x)$ such that $g(\cdot)$ is a deterministic function, and $f(\cdot)$ is a surjective function. Therefore, $p(X=f(y)|Y=y)=1$, or equivalently, $p_{X|Y}(x|y)=\delta(x-f(y))$. Given a particular input $Y=y$, a restoration algorithm produces an estimate $\hat{X}$ according to some distribution $p_{\hat{X}|Y}(\cdot|y)$ such that the estimate $\hat{X}$ is statistically independent of $X$ given $Y$ . For deterministic algorithms, $p_{\hat{X}|Y}(\cdot|y)$ is a delta function for every $y$, while for stochastic algorithms it is a non-degenerate distribution. Recall,
(53) $H(\theta|X)=\sum_{x}p(x)H(\theta|X=x)=-\sum_{x}p(x)\sum_{\theta}p(\theta|x)\log p(\theta|x)=-\sum_{x,\theta}p(x,\theta)\log p(\theta|x).$
Assuming $\theta-Y-X$ forms a Markov chain, $P(\theta,X)=P(\theta,X|Y=T(X))$, and $H(\theta|X)\leq H(\theta|Y)$, with equality if $g(\cdot)$ is a bijective function. Thus, in case that $\mathcal{F}:\mathcal{Y}\rightarrow\mathcal{X}$ is a surjective mapping fully captured by the estimator of $\hat{X}$, i.e., $p_{X|Y}(\cdot|y)=p_{\hat{X}|Y}(\cdot|y)$, as stated in theorems 5.3-5.5 an algorithm that is able to extract the source signal conditional distribution would yield a parameter estimation result that is as accurate as the actual source and as the measurement. The simplest example would be $Y=aX$, where $a>0$ is a constant.
In the case that $y=g(x)$ where $g:\mathcal{X}^{n}\rightarrow\mathcal{Y}^{d}$ is a non-invertible deterministic degradation such that $p_{Y|X}(y|x)=\delta(y-g(x))$, Ohayon et al., 2023 [Ohayon:2023a] proved that when a restoration algorithm enforces both perfect perceptual quality , i.e., $p_{\hat{X}}=p_{X}$ and perfect data fidelity, that is y=g(x^)y=g($\hat${x}) (equivalent to $p_{Y|\hat{X}}=p_{Y|X}$), it must be a sampler from the posterior distribution $p_{X|Y}$, and is thus required to be stochastic. In other words, these are ill-posed inverse problems in which $X$ cannot be retrieved from $Y$ with zero error ($p_{X|Y}(\cdot|y)$ is not a delta function for almost every $y$). Problems such as image colorization, inpainting, single-image super-resolution, JPEG-deblocking, and more all follow this assumed structure. Thus, in these cases, as indicated in Theorems 5.3-5.7, in order to maintain the same level of accuracy for parameter estimation after the reconstruction, there is need for prior knowledge in the restoration process, such as an oracle of $\theta$, and/or a restoration process that does not depend on $\theta$.
@@ 6.2 When $\hat{\theta}_{Y}$ and $\hat{\theta}_{\hat{X}}$ coincide
Let us assume, $Y=X+Z$, where $Z\sim\mathcal{N}(0,\sigma_{\mathrm{n}}^{2})$ and $X$ are independent. Hence $Z$ and $\theta$ are independent. That is, $\theta-X-Y$ forms a Markov chain. As known, generally speaking, in this case, averaging examples of $Y$ yields $X$, that is, $E[Y]=X$.
@@
Now, assume $Y=X+Z$, then $Y\sim\mathcal{N}(\mu,\sigma^{2}_{\mathrm{y}})$ such that $\sigma^{2}_{\mathrm{y}}=\sigma^{2}_{\mathrm{x}}+\sigma^{2}_{\mathrm{z}}$. Assuming that $\theta=\mu\sim\mathcal{N}(\mu_{\mathrm{m}},\sigma^{2}_{\mathrm{m}})$, we have $I(\theta,X)=\frac{1}{2}\mathrm{log}(1+\frac{\sigma^{2}_{\mathrm{m}}}{\sigma^{2}_{\mathrm{x}}-\sigma^{2}_{\mathrm{m}}})$, and $I(\theta,Y)$$\frac{1}{2}\mathrm{log}(1+\frac{\sigma^{2}_{\mathrm{m}}}{\sigma^{2}_{\mathrm{y}}-\sigma^{2}_{\mathrm{m}}})$. As well as $J^{-1}_{m,X}(\theta)=\frac{\sigma_{\mathrm{x}}^{2}}{m}$ and $J^{-1}_{m,Y}(\theta)=\frac{\sigma_{\mathrm{y}}^{2}}{m}$. At first glance, this may seem as a contradiction, since, obviously $I(\theta,X)>I(\theta,Y)$. Yet, provided that we would be able to recover $\hat{x}$, such that $p_{\hat{X}|\theta}(x|\theta)=p_{X|\theta}(x|\theta)$, which may be approximately possible in practice for some cases, under sufficiently low noise levels, we should be able to achieve the minimal CRB. Recall, that for a memory-less source we have
(55) $I(\mathbf{x},\mathbf{y})\geq\sum^{m}_{i=1}I(\mathbf{x}_{i},\mathbf{y}_{i}),$
@@
An optimal linear estimator for $X$ in the above example would be $\hat{x}=ay+b$ such that $a=\frac{\sigma_{x}}{\sigma_{y}}$, $b=E[Y](1-\frac{\sigma_{\mathrm{x}}}{\sigma_{\mathrm{y}}})$. In that case, in accordance with Theorem 5.5, clearly $P_{\hat{X}}(x)=P_{X}(x)$, yet obviously $\forall\theta\ :\ \hat{x}\neq x$.
@@ A Gaussian encoder (variance preserving) at time step $t$,
(56) $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\sqrt{\alpha_{t}}\mathbf{x}_{t-1},(1-\sqrt{\alpha_{t}})\mathbf{I}).$
We are interested in learning conditionals $p_{\phi}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, so that we can simulate new data. Intuitively, the forward process gradually adds noise to the observation $\mathbf{x}_{0}$, whereas the generative process gradually denoises a noisy observation. If we assume $\theta=\mathbf{x}_{0}$ is the clean image, and $\mathbf{x}=\mathbf{x}_{t}$ such that $0<t<T$ and $\mathbf{y}=\mathbf{x}_{T}$ is the noise image. The generative process approximates an intractable reverse process $p_{\phi}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$. The training loss [Ho:2020] uses KL divergence to directly compare $p_{\phi}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ against forward process posteriors, which are tractable when conditioned on $\mathbf{x}_{0}$.
@@ 6.3 Sparse Inversion
(57) $=\arg\min_{\mathbf{x}\in\mathbb{R}^{N}}\quad\log{p_{Y|X}(\mathbf{y}|\mathbf{x})}-\log{p_{X}(\mathbf{x})}.$
@@
(58) $\hat{\mathbf{x}}=\arg\min_{\mathbf{x}\in\mathbb{R}^{N}}\quad\frac{1}{2\sigma_{\mathrm{z}}^{2}}\|\mathbf{y-Hx}\|_{2}^{2}-\lambda\|\mathbf{x}\|_{1},$
which is considered in the literature as the well known sparse coding $\ell_{1}$ optimization problem (see [Elad:2010], Chapter 6). The first term is (58) is the data fidelity term, and the second term is the regularization term. The sparse coding recovery estimation is an extensively studied problem [Granda:2013, Granda:2013B, Pereg:2017, Pereg:2021]. In [Pereg:2019] we proved that the bound on the estimation of $\mathbf{x}$ under sufficient separation between adjacent non-zero values (spikes) depends on the noise level. In the absence of noise, the MAP estimation should be perfect. However, note that solving (58) accurately requires the prior knowledge of $\lambda$.
@@
(59) $s^{m}(\lambda)=\frac{\partial}{\partial\lambda}\log p^{m}_{\lambda}(\mathbf{x})=\frac{\partial}{\partial\lambda}\big(m\log(\lambda)-\lambda\sum_{i=1}^{m}\|\mathbf{x}_{i}\|_{1}\big).$
@@
(60) $J^{m}_{X}(\lambda)=mJ_{X}(\lambda)=-mE\Bigg\{\Bigg(\frac{\partial^{2}}{\partial\lambda^{2}}\log p_{\lambda}(\mathbf{x})\Bigg)^{2}\Bigg|\lambda\Bigg\}=\frac{m}{\lambda^{2}},$
In other words, hypothetically, if we were to first perfectly denoise the signal such that $\|\hat{\mathbf{x}}\|_{1}=\|\mathbf{x}\|_{1}$, assuming that our denoising mechanism is errorless, we have $\lambda^{-1}_{\hat{X}}=\|\hat{\mathbf{x}}\|_{1}$ ($m=1$). That said, by Theorem 5.3 we know that even in that case $\Delta_{Y}=\Delta_{\hat{X}}$. In other words, if we were to estimate $\lambda$ directly from the measurement, then our expected bound on the parameter estimation error is the same. If we solve the MAP for noisy data the upper-bound for $\hat{\mathbf{x}}$ is given by eq. (78),(82) in Appendix C, and it cannot be guaranteed that $\lambda_{\hat{X}}=\lambda$, namely, $\Delta_{Y}\leq\Delta_{\hat{X}}$. Interestingly, in the noisy case, denoising the measurement will not improve the following expected estimation error.
@@
[…]
7 Discussion &\& Conclusions
When working on inverse problems in different domains, we tend to assume that the inversion process cleans the data and extracts the relevant information. We often forget that the process could also be inherently biased and inherently inclined to lose information. This study broadly analyzes parameter estimation and domain shift in inverse problems. We introduced general definitions for perfect perceptual estimators in the context of parameter estimation in continuous and discrete settings, and presented theoretical bounds for reliable prediction. Observing that domain randomization and data augmentation methods attempt the build models that generalize over several distributions, we highlight a profound vulnerability of this approach in the Double Meaning Theorem (Theorem 3.2), which often leads to mode collapse in signal restoration tasks. Our theoretical findings are confirmed in experimental examples. Future work can analyze other specific distributions depending on parameter estimation, such as: Bernoulli distribution, and mixture models e.g., product of experts [Hinton:1999], experimentally investigate other types of domain changes and adversarial attacks in different applications, and develop coping strategies. Our future work will investigate how bias during training affects the classification decision, and the effect of number of measurements, and sample size.
@@ Appendix A Information-Theory Preliminaries
A.1 Ergodic Process
Generally speaking, when a signal is a function of time (or space), an ergodic signal is a stationary signal whose time estimated moments approximate its statistical ensemble moments.
Definition A.1 (Mean-Ergodic Processes [Papoulis:1991]).
Considering a continuous signal $x(t)$, where the index $t$ denotes time. The signal mean is $\eta=E[x(t)]$ and its average is $\eta_{T}=\frac{1}{2T}\int_{-T}^{T}x(t)dt$. A mean-ergodic process obeys
(61) $\lim_{T\rightarrow\infty}\eta_{T}=\lim_{T\rightarrow\infty}\frac{1}{2T}\int_{-T}^{T}x(t)dt=E[x(t)]=\eta.$
Definition A.2 (Autocorrelation-Ergodic Processes [Papoulis:1991]).
A random stationary process $x(t)$ is ergodic in autocorrelation if the time-averaged autocorrelation of a single realization equals the ensemble autocorrelation,
(62) $\lim_{T\rightarrow\infty}\mathcal{R}_{T}(\tau)=\lim_{T\rightarrow\infty}\frac{1}{2T}\int_{-T}^{T}x(t)x(t+\tau)dt=E[x(t)x(t+\tau)]=\mathcal{R}(\tau).$
A.2 Efficiency
@@
(63) $V(\hat{\theta})=\frac{J^{-1}(\theta)}{\mathrm{Var}(\hat{\theta})}$,
@@
(64) $\mathrm{Var}(\hat{\theta})=E(\theta-\hat{\theta})^{2}.$
@@ A.3 Consistency
A.4 Minimal Sufficient Statistic
Definition A.3 (sufficient statistic [ThomasCover:2006]).